OpenStudy (anonymous):

A hiking club has 9 members. How many ways can they choose a president, vice president, secretary, and treasurer from among the members? A. 6561 B. 36 C. 126 D. 3024

OpenStudy (anonymous):

so there are 4 spots, right?

OpenStudy (anonymous):

one of the 9 members will be president. how many does that leave?

OpenStudy (anonymous):

8, right? 9-1=8. after that 8-1 =7. 7-1=6. 9x8x7x6=3024

OpenStudy (anonymous):

so D) 3024.

OpenStudy (anonymous):

Excellent explanation thank you very much!!!

OpenStudy (anonymous):

no problem